OpenStudy (mousejkl):

Which is an equation in point-slope form of the line that passes through the points  (4, 5) (4, 5) and (−3, −1) (−3, −1) ? y+1= 6 7 (x+3) y+1=67(x+3)

OpenStudy (mousejkl):

anyone

OpenStudy (mousejkl):

this

OpenStudy (will.h):

Thanks for posting a screenshot that helps more So we have the pointS (4,5) (-3,-1) Can you find the slope? Slope = Y2-Y1/X2-X1

OpenStudy (mousejkl):

good what elels

OpenStudy (will.h):

Find the slope 1st. show me ur work

OpenStudy (will.h):

Find the slope and then use the following formula Y - Y1 = m(x - x1) Where y1= -1 and x1= -3 And m is the slope Hope that helps

OpenStudy (mathmale):

Could you possibly make your post clearer? Why the duplications in: (4, 5) (4, 5) and (−3, −1) (−3, −1) ??? See and follow the advice from Will.H.

OpenStudy (mousejkl):

helllllllllllllllllllllllllllllllllllllllllllllllllllllllllllllllllllp

OpenStudy (mathmale):

Is this a problem separate from the at the top of this page? If so, what are the instructions for this 2nd problem? Are y ou able to identify the coordinates of the point at which the 2 lines intersect?